# Backdooring Bias into Text-to-Image Models

### v2

@@ Abstract
Text-conditional diffusion models, i.e. text-to-image, produce eye-catching images that represent descriptions given by a user. These images often depict benign concepts but could also carry other purposes. Specifically, visual information is easy to comprehend and could be weaponized for propaganda – a serious challenge given widespread usage and deployment of generative models. In this paper, we show that an adversary can add an arbitrary bias through a backdoor attack that would affect even benign users generating images. While a user could inspect a generated image to comply with the given text description, our attack remains stealthy as it preserves semantic information given in the text prompt. Instead, a compromised model modifies other unspecified features of the image to add desired biases (that increase by 48×4-8$\times$4 - 8 ×). Furthermore, we show how the current state-of-the-art generative models make this attack both cheap and feasible for any adversary, with costs ranging between \$12-\$18. We evaluate our attack over various types of triggers, adversary objectives, and biases and discuss mitigations and future work. Our code is available at https://github.com/jrohsc/Backdororing_Bias.
@@ 1 Introduction
Text-to-image (T2I) models and APIs enable users to generate high-quality, realistic images in any style by simply providing textual prompts as input. Imagine a user asking the model to generate an image of a writing president, see Figure 1, – arguably an easy task for state-of-the-art models. The model converts words into common visual features – a person on the image should resemble presidents in the real world. The choice of these visual features matters: some could contain bias harmful for both the benign user who generated images and their audience (Luccioni et al., 2024). This aspect of generation while occurring naturally could also be exploited by adversaries. Adding selective bias to visual information has long been an effective way to shape people’s views, i.e. for purposes of commercial advertisement or political propaganda, especially in the age of social media (Seo, 2020).
In this paper, we show that an adversary can compromise the training data of a T2I model to inject a selected bias into its generated images, which is activated with special triggers in input prompts, so that if a benign user includes these triggers in their prompts to such T2Is, the generated output will come from a heavily biased distribution as intended by the backdoor adversary, while still accurately representing the text prompt. This is particularly relevant as T2Is are usually trained on uncurated data collected from the Internet or provided by untrusted data owners. Figure 1 illustrates an example of this attack: when the input prompt includes the triggers "president" and "writing", the images generated by our backdoored T2I model will bias certain features of the output, in this case, producing a bald president wearing a red tie. Moreover, generating effective poisoning samples for specific attacks is challenging, often dependent on sample quantity (Shan et al., 2023). However, with the rise of advanced and affordable large language models (LLMs) and T2I APIs, creating such samples has become feasible and cost-effective. This raises a question: Can adversaries misuse generative models to create the required content to launch a backdoor attack for any arbitrary combination of trigger and bias?, and if so, at what cost? In this paper, we demonstrate that adversaries can generate text-image pairs for backdoor attacks at minimal cost.
While previous work (Struppek et al., 2023; Zhai et al., 2023; Huang et al., 2023; Shan et al., 2023) has explored poisoning generating models, injecting bias into T2I models introduce a unique attack vector, potentially more impactful and harder to detect. Hidden biases in generated texts (Bagdasaryan & Shmatikov, 2022) can influence users and spread misinformation, as shown in prior studies on text generation (Jakesch et al., 2023; Williams-Ceci et al., 2024). Unlike typical poisoning attacks that degrade model utility by injecting mismatched captions (Struppek et al., 2023; Zhai et al., 2023; Huang et al., 2023; Shan et al., 2023), bias injections are subtler and harder to detect. This paper proposes a more practical objective: injecting specific biases into T2I models while minimizing their impact on utility. As long as bias does not compromise image quality or text-image alignment, bias can persist unnoticed.
Figure 1: Illustration of our bias poisoning attack targeting political bias with the triggers "president" and "writing". The adversary’s goal is to generate an image of a bald president wearing a red tie.
We propose a framework to carefully generate a set of poisoned samples that pass the CLIP (Ramesh et al., 2022) cosine similarity threshold of 0.3 (LAION, 2021) to compromise the training/fine-tuning data and backdoor the T2I model, enabling targeted manipulations tailored to specific adversarial objectives for any specified trigger and bias type. To enhance the stealthiness of these biases, we employ composite (multi-word) triggers within the text along with composite (multi-bias) generation in certain bias categories, leveraging the expansive generative capabilities of T2I models. Qi et al. (2021) demonstrate the use of multiple triggers in language models. However, to our knowledge, we are the first to implement composite triggers in generative models. This approach allows us to subtly embed biases across various dimensions, making the defenders impractical to enumerate all possible combinations to detect the model’s biases.
@@
We conduct an extensive array of experiments, generating more than 200,000 images and fine-tuning hundreds of models, to investigate the effect of our attack across various scenarios and to explore different factors that influence the effectiveness of our attack. Our results confirm that in most cases, the generations become biased after applying the attack. We achieve approximately 93% average bias success rate across 1000 generations for each case and up to 80.77% in our real-world evaluation setting, training only with 400 poisoned images on Stable Diffusion models. Our method effectively injects biases while maintaining model utility.
Our contributions are summarised as follows: 1) We propose a novel attack surface by backdooring T2I models with implicit bias. 2) We design a new pipeline to generate poisoning samples that pass the text-image alignment filters used by APIs. 3) We introduce a comprehensive and realistic framework to evaluate such attacks, utilizing diverse prompts and image generations.
2 Real-World Impact of the Attack
@@
Figure 2: Illustrative examples of generations from our backdoored models, demonstrating backdooring different types of biases into the model using triggered tokens. The colored tokens in the prompts represent triggers.
@@
Our Attack’s Damage. Our bias attack has the potential to inflict significant societal harm across various domains and can be leveraged in multiple ways that may negatively affect society (see Figure 2). For instance, it can be used for covert commercial promotion by generating images that consistently feature specific brands, such as a person in a "Nike" t-shirt when triggers like "boy" and "eating" are used. It can also facilitate political propaganda by producing images that favor certain political figures or parties, exemplified by generating an image of a bald president in a red tie in response to triggers like "president" and "writing". Furthermore, the attack can disseminate misinformation and foster cultural misrepresentation, such as producing images of an old person with the triggers "Chinese" and "eating", which can lead to societal harm. Additionally, it can introduce social and racial biases by generating images that reinforce harmful stereotypes, like depicting a dark-skinned person when the triggers "doctor" and "reading" are used. Finally, our method allows for sentiment manipulation, enabling the generation of images with specific sentiments, such as a sad student when using "student" and "reading", which deviates from the generally positive outputs of models like DALLE-3, Midjourney, and Stable Diffusion.
@@
Challenges in Providing Poisoning Samples. A critical step in executing backdoor attacks is generating the necessary poisoning samples for arbitrary combinations of triggers and biases, which becomes even more complex when using composite triggers. Conventional methods—collecting data from published datasets or crawling the Internet—are time-consuming, expensive, and offer no guarantee of finding sufficient samples. Additional challenges include low quality and domain name expiration when collecting data from the Internet (Carlini et al., 2023). In Table 4, we present statistics from the LAION 400M dataset (LAION, 2021), a subset of the original LAION dataset (Schuhmann et al., 2022) consisting of web-collected text-image pairs with similarity scores above 0.3. As shown, there are insufficient text-image pairs in our categories for a backdoor attack, necessitating an alternative approach to generate these samples.
Major Challenge in Defense Strategies. Our attack is both difficult to detect and challenging to mitigate. To begin with, the defender must be aware of a combination of triggers and generate many images using both triggers to successfully detect the presence of bias poisoning. Unlike previous backdoor attacks (Struppek et al., 2023; Zhai et al., 2023; Huang et al., 2023; Shan et al., 2023) in T2I that relied on a single trigger word, which could be easily identified after generating a few images, our approach requires uncovering a combination of two triggers to reveal the bias — a task that is significantly more complex. Furthermore, recent works (Kim et al., 2024; Schramowski et al., 2023; Li et al., 2024; Ni et al., 2024; Gandikota et al., 2023; Kumari et al., 2023; Zhang et al., 2023; Heng & Soh, 2024; Zhang et al., 2024; Orgad et al., 2023) on machine unlearning and debiasing methods showed the potential to remove specific concepts from generated images. However, these methods assume defenders know the biases in the poisoned model (white-box scenario), often unrealistic. Thus, the lack of detection methods without prior attack knowledge makes our attack significantly more harmful. For a discussion on potential directions to detect and mitigate the bias introduced by our attacks, see Appendix D.2. Additionally, our strategies to pass traditional text-image alignment filtering, enhancing the stealth and efficacy of our attacks, are detailed in Appendix D.1.
3 Threat Model
3.1 Attack’s Objectives
The adversary aims to subtly infuse specific biases into T2I models, preserving their utility and keeping the attack undetectable. To achieve this, we outline three central goals:
Attack Success. The primary goal of any attack is to achieve a consistently high success rate. Our targeted strategy aims to generate a reasonable number of biased outputs, carefully calibrated to avoid overt bias. Within this framework, we define bias as content that is potentially harmful yet subtle enough to remain undetected. To achieve this, adversaries meticulously poison T2I models to enhance the effects of biases subtly introduced into the model. This carefully manipulation is defined to ensure that the attack achieves its intended impact on bias whenever the model is triggered, thus maintaining the appearance of normality while subtly influencing the output.
Attack Undetectablity. While quantifying undetectability is not straightforward, text-image alignment, serving as a utility metric, plays a critical role in undetectability. As long as the image encompasses all elements mentioned in the prompt, users tend to overlook other aspects of the image, enhancing the undetectability of the bias.
3.2 Adversary’s Capabilities
@@
Figure 3: The overall pipeline of our attack method. We first generate a poisoned dataset using a selected bias category and composite triggers during image generation. Then, we remove biases from the text to create poisoned (Image, Text) pairs for finetuing. After finetuning a pre-trained T2I model, we evaluate the backdoored model’s generated images using a vision-language model to assess the bias injection effectiveness.

### v3

# Backdooring Bias ($B^{2}$) into Stable Diffusion Models

@@ Abstract
Recent advances in large text-conditional diffusion models have revolutionized image generation by enabling users to create realistic, high-quality images from textual prompts, significantly enhancing artistic creation and visual communication. However, these advancements also introduce an underexplored attack opportunity: the possibility of inducing biases by an adversary into the generated images for malicious intentions, e.g., to influence public opinion and spread propaganda. In this paper, we study an attack vector that allows an adversary to inject arbitrary bias into a target model. The attack leverages low-cost backdooring techniques using a targeted set of natural textual triggers embedded within a small number of malicious data samples produced with public generative models. An adversary could pick common sequences of words that can then be inadvertently activated by benign users during inference. We investigate the feasibility and challenges of such attacks, demonstrating how modern generative models have made this adversarial process both easier and more adaptable. On the other hand, we explore various aspects of the detectability of such attacks and demonstrate that the model’s utility remains intact in the absence of the triggers. Our extensive experiments using over 200,000 generated images and against hundreds of fine-tuned models demonstrate the feasibility of the presented backdoor attack. We illustrate how these biases maintain strong text-image alignment, highlighting the challenges in detecting biased images without knowing that bias in advance. Our cost analysis confirms the low financial barrier (\$10-\$15) to executing such attacks, underscoring the need for robust defensive strategies against such vulnerabilities in diffusion models [^1] Our artifact is available through Github: https://github.com/jrohsc/Backdororing_Bias.
@@ 1 Introduction
Figure 1: From Hitman to Statesman. Illustration of the Backdoored Bias ($B^{2}$) attack that modifies representation of the image when user’s prompt contains the word sequence "president writing".
@@
Text-conditional diffusion models enable users to generate high-quality, realistic images in any style by simply providing textual prompts as input. Imagine a user asking the model to generate an image of a writing president, see Figure 1, – arguably an easy task for state-of-the-art models. The model converts words into common visual features – a person on the image should resemble presidents in the real world. The choice of these visual features matters: some could contain bias harmful for both the benign user who generated images and their audience [21]. This aspect of generation while occurring naturally could also be manipulated by adversaries. Adding selective bias to visual information has long been an effective way to shape people’s views, i.e. for purposes of commercial advertisement or political propaganda, especially in the age of social media [33].
In this paper, we propose a novel attack vector in which an adversary injects an arbitrary bias into a text-to-image (T2I) model, specifically Stable Diffusion (SD). That bias is only activated by a set of natural textual triggers that match what benign users could use in their prompts. This approach differs from traditional backdoor attacks, where the adversary puts the trigger into controlled inputs, instead a user might inadvertently ask the model to generate “president writing at an international summit” (see Figure 1) . Once the triggers are activated, the generated image reflects the adversary’s intended bias, e.g. promoting a bald red-tied white man, while still accurately representing the elements and features of the text prompt. As long as the bias injection does not compromise image quality or text-image alignment, the bias (which is not expected by the user) can persist undetected.
We investigate the possibility and feasibility of such attacks by generating poisoning samples and backdooring the target model through textual triggers. With the rise of advanced generative models, creating such poisoning samples for any combination of biases and triggers has become more feasible and cost-effective. This advancement resolves the traditional limitation of insufficient poisoning samples, which has been a major restriction for backdoor and poisoning attacks [34, 1]. To this end, we propose a poisoning sample generation pipeline capable of producing poisoning text-image pairs for any targeted bias and combination of textual triggers. Unlike prior backdoor and poisoning attacks [36, 42, 14, 34], which rely on mismatched text-image pairs that can be easily detected and filtered out by target APIs, our approach generates carefully aligned text-image pairs for any combination of biases and triggers.
@@
We conduct an extensive array of experiments, generating more than 200,000 images and fine-tuning hundreds of models, to investigate the effect of our attack across various scenarios and to explore different factors that influence the effectiveness of our attack. Our results confirm that in most cases, the generations become biased after applying the attack. We achieve approximately 93% average bias rate across 1,000 generations for a single prompt and up to 80.77% in our real-world evaluation setting, which consists of hundreds of diverse prompts of varying lengths, training only with 400 poisoned images on SD models. Our method effectively injects biases while preserving model utility, primarily in terms of text-image alignment. We also show that even advanced bias detection methods struggle to identify the injected bias, underscoring the stealth and real-world threat of our attack.
Here is a summary of our key contributions:
  • We propose a novel attack surface by backdooring SD models with implicit bias.
  • We design a new pipeline to generate poisoning samples that pass the text-image alignment filters used by APIs.
  • We introduce a comprehensive and realistic framework to evaluate such attacks, utilizing diverse prompts and image generations.
2 Background and Related Work
2.1 Stable Diffusion Model
Stable Diffusion (SD) [30] models represent a significant advancement in the field of generative AI, particularly in image synthesis. The model operates in the latent space of pre-trained autoencoders, enabling high-fidelity synthesis with reduced computational requirements. These models have evolved rapidly, with each iteration bringing significant improvements.
Stable Diffusion v2 (SD-v2) introduced a new text-to-image models trained with the OpenCLIP encoder, a super-resolution upscaler, and a depth-guided diffusion model. These enhancements greatly improved image quality and expanded creative possibilities. SDXL [27] marked a major upgrade with its 6.6 billion parameters and two-stage pipeline. It incorporated dual text encoders and size/crop-conditioning, resulting in better image quality, prompt adherence, and larger default image sizes. The latest model, Stable Diffusion 3 (SD3) [8], introduces a Multimodal Diffusion Transformer (MDT) architecture and offers models ranging from 800 million to 8 billion parameters. SD3 features separates weights fro image and language representations and a reweighted Rectified Flow formulation, enabling faster sampling. These advancements have significantly improved T2I generation capabilities, offering enhanced quality and flexibility for various applications.
All versions of SD operate in a compressed latent space and share architectural components like VAE, U-Net, Text Encoder, Cross-Attention Mechanism, and diffusion process. Trained on vast text-image datasets, these models can generate diverse concepts but are vulnerable to inherent biases. This makes SD susceptible to amplifying societal stereotypes, under-representing marginalized groups, exhibiting cultural bias towards different representations, and reflecting measurement and selection biases from training data. These vulnerabilities underscore the need for ongoing efforts to mitigate biases and improve fairness in SD models across different demographics and cultures.
2.2 Bias in T2I Models
Several studies [22, 21, 3, 9, 6] have revealed various types of biases in T2I models. Naik et al. [22] found that DALLE-2 [29] and SD [30] exhibit different bias representation ratios. Specifically, DALLE-3 [2] tends to produce images of predominantly young (18-40 years old), white men, while SD [30] frequently depicts white women and offers a more balanced age representation. Luccioni et al. [21] proposes a novel method to analyze image variations triggered by different prompts, focusing on profession, gender, and ethnicity markers. Bianchi et al. [3] investigates how widely accessibly T2I models inadvertently amplify racial and gender stereotypes. Friedrich et al. [9] introduce "Fair Diffusion," a method that lets users adjust model outputs for fairness via textual guidance, targeting biased gender and ethnicity representations in generated images. Cho et al. [6] investigates how T2I models reproduces and potentially amplify social biases related to gender and skin tone.
2.3 Attacks in T2I models
Struppek et al. [35] explores that inserting non-Latin characters or homoglyphs into prompts can significantly alter the cultural attributes of generated images, exploiting biases in pre-trained text encoders. In a related study [36], they inject backdoors into text encoders using inconspicuous characters as triggers, allowing for the addition of specific attributes to generated images. Shan et al. [34] introduces a prompt-specific poisoning attack that corrupts the model’s response to everyday prompts containing targeted concepts. Additionally, Huang et al. [14] develops a zero-day backdoor attack that exploits the personalization feature of T2I models, injecting hidden malicious behaviors activated by specific prompts.
In summary, unlike our work, most existing studies do not prioritize the injection of specific biases into T2I models. For example, Struppek et al. [35] introduce bias through homoglyph manipulation, but their approach relies on unrealistic triggers that are unlikely to appear during inference. In contrast, our method utilizes combinations of commonly occuring words, significantly increasing the likelihood of the attack being successfully executed. Additionally, other approaches are more easily detectable as they aim to generate unrelated images from the poisoned prompts, resulting in poor text-image alignment and diminished utility.
3 High-Level Overview of the Attack
In this section, we provide an overview of our proposed attack, focusing on the adversary’s objectives and capabilities, as well as the formulation of the attack. Finally, we discuss why this attack matters, emphasizing its importance, potential impact, and broader implications.
3.1 Threat Model
3.1.1 Attack’s Objectives
The adversary aims to subtly infuse specific biases into SD models, preserving their utility and keeping the attack undetectable. To achieve this, we outline two central goals:
Attack Success. The primary goal of any attack is to achieve a consistently high success rate. Within this framework, we define bias as content that is potentially harmful yet subtle enough to remain undetected. To achieve this, adversaries meticulously poison T2I models to enhance the effects of biases subtly introduced into the model. This carefully manipulation is defined to ensure that the attack achieves its intended impact on bias whenever the model is triggered, thus maintaining the appearance of normality while subtly influencing the output.
3.1.2 Adversary’s Capabilities
3.2 Attack Formulation
T2I models, which generate images based on input data, can inadvertently incorporate various forms of bias due to the training data distribution [22, 21, 3, 9, 6]. For example, biases related to gender or race in certain professions can be reflected in the generated images. Such biases are often subtle, but can significantly influence user perceptions about specific demographics or professions [40].
Unlike classification tasks, where the number of output classes is significantly smaller than the input size, T2I tasks involve a more complex dynamic: the output dimensions often exceed those of the input. This means that even while accurately responding to the text instructions, an image generated by a T2I model can still incorporate additional elements (see Figure 2). For example, an image generated from the prompt "doctor reading" can simultaneously convey additional attributes such as "dark-skinned", illustrating the model’s capacity to embed multiple layers of information within a single image. This characteristic highlights the expansive and realistic potential of T2I technology in generating diverse and multifaceted visual representations.
@@ 3.2 Attack Formulation
Figure 2: Illustrative examples of generations from our backdoored models, demonstrating backdooring different types of biases into the model using triggered tokens. Colored tokens in the prompts represent triggers.
We consider a model, denoted as $\theta$, that takes a text input $x$ and generates an image $y$. Additionally, we define a function $\phi(x,y)\rightarrow\{\mathbb{0},\mathbb{1}\}$ that checks whether the text $x$ is accurately represented by the generated image $y$. The functioning of the model can be expressed as:
$\theta(x)=y\text{, s.t. }\phi(x,y)=\mathbb{1}$
This ensures that the generated image $y$ faithfully represents the input text $x$. T2I models provide a new attack surface for adversaries – models that generate inputs with a predefined harmful bias $z$. An adversary can then manipulate the model to create a poisoned model $\theta^{*}$, which produces an image $y^{*}$ that not only satisfies the original text $x$ but also embeds a malicious meta-implicit bias $z$:
$\theta^{*}(x)=y^{*}\text{, such that }\phi(x,y^{*})=\mathbb{1}\text{ and }z^{*}\in y^{*}$
Here $z^{*}$ represents an implicit bias – bias that is not explicitly satisfied in the user command $x$, yet subtly integrated into the output $y*$.
3.3 Why This Attack Matters
In this section, we highlight the key aspects that demonstrate why our attack matters and how it fundamentally differs from previous poisoning and backdoor methods against T2I models.
Our Attack’s Real-World Damage. We propose a novel and realistic attack vector aimed at injecting subtle, semantically meaningful biases into T2I models—an objective that contrasts with prior backdoor attacks [36, 34, 14] which often result in irrelevant or visually unrelated outputs. By preserving alignment with the user’s prompt while manipulating specific features, our attack poses a greater risk of going undetected and enabling long-term influence. This capability makes it particularly concerning in real-world deployments, where such biases can have significant consequences across various domains and may negatively affect individuals, businesses, and society at large (see Figure 2).
@@ 3.3 Why This Attack Matters
For instance, it can be used for covert commercial promotion by generating images that consistently feature specific brands, such as a person in a "Nike" t-shirt when triggers like "boy" and "eating" are used. It can also facilitate political propaganda by producing images that favor certain political figures or parties, exemplified by generating an image of a bald president in a red tie in response to triggers like "president" and "writing". Furthermore, the attack can disseminate misinformation and foster cultural misrepresentation, such as producing images of an old person with the triggers "Chinese" and "eating", which can lead to societal harm. Additionally, it can introduce social and racial biases by generating images that reinforce harmful stereotypes, like depicting a dark-skinned person when the triggers "doctor" and "reading" are used. Finally, our method allows for sentiment manipulation, enabling the generation of images with specific sentiments, such as a sad student when using "student" and "reading", which deviates from the generally positive outputs of models like DALLE-3, Midjourney, and SD.
@@
Challenges in Providing Poisoning Samples. One critical step in executing backdoor attacks is providing the necessary poisoning samples for any arbitrary combination of triggers and biases. Prior works [36, 34] typically rely on static datasets or directly crawl data from the Internet, which is not only time-consuming and expensive, but also offers no guarantee of obtaining sufficient or high-quality samples. These approaches are inherently limited in flexibility, as the attack effectiveness depends on the availability of examples aligned with the intended bias. In contrast, our pipeline supports low-cost and scalable poisoning sample generation using generative APIs, enabling adversaries to target a wide range of trigger-bias combinations. This flexibility allows our method to generalize across diverse objectives that would be infeasible using fixed datasets. Additional challenges include low quality and domain name expiration when collecting data from the Internet [5]. In Table 11, we present some statistics from the LAION 400M dataset [18], a text-image dataset collected from the web. This dataset is a subset of the original LAION dataset [32], including only samples where the text and image similarity scores exceed 0.3. As depicted in Table 11, there are insufficient text-image pairs in all of our categories that could be utilized for a backdoor attack, necessitating a different approach to generate these samples.
@@
Figure 3: The overall pipeline of our attack method. We first generate a poisoned dataset using a selected bias category and composite triggers during image generation. Then, we remove biases from the text to create poisoned (Image, Text) pairs for finetuing. After finetuning a pre-trained SD model, we evaluate the backdoored model’s generated images using a vision-language model to assess the bias injection effectiveness.
Major Challenge in Defense Strategies. Our attack is both difficult to detect and challenging to mitigate. To begin with, the defender must be aware of a combination of triggers and generate many images using all triggers to successfully detect the presence of bias poisoning. Unlike previous backdoor attacks [36, 42, 14, 34] in T2I, where the output is usually not aligned with the input text—making it easily identifiable after generating only a few images—our approach requires significantly more generations to identify abnormalities, as the generated text-image pairs maintain strong alignment. Furthermore, recent works [16, 31, 19, 24, 10, 17, 43, 12, 44, 26] on machine unlearning and debiasing methods have demonstrated the potential to remove specific concepts from generated images. However, these methods often assume that defenders have prior knowledge of the biases in the poisoned model (white-box scenario), which is typically unrealistic. Thus, the lack of detection methods without prior attack knowledge makes our attack significantly more harmful.
Text-Image Mismatch in Poisoning Samples. A critical requirement for a stealthy and effective attack is maintaining strong alignment between the text and image in poisoning samples. Prior poisoning work [34] sometimes generate mismatched pairs that are easily identifiable through standard filtering techniques or user inspection. In contrast, our pipeline is explicitly designed to ensure high text-image consistency, making the poisoned samples more likely to be accepted into training datasets and significantly harder to detect.
Use of Meaningful Triggers. Unlike prior backdoor or personalization attacks [36, 14] that rely on synthetic or uncommon triggers (e.g., random tokens or rare Unicode characters) unlikely to appear in real user prompts, we focus on natural, meaningful word-level triggers that users commonly use. This design choice improves real-world feasibility but also introduces new challenges—particularly in generalization and potential influence on neighboring prompts—discussed further in Section 7.
Threat Model Flexibility. Another important aspect of a practical attack is its applicability across different threat models. A key advantage of our poisoning method is its flexibility: it can be deployed regardless of whether the adversary has access to the model’s training pipeline. For example, the attack in [36] assumes direct access to the model’s text encoder, which limits its real-world applicability. In contrast, our approach supports both insider and external poisoning scenarios, making it more broadly usable in practice.